OpenStudy (anonymous):

Simplify 6x^0/y^−5 .

OpenStudy (anonymous):

Here's my guess >> \[x^0 = 1\] So \[6x^0 = 6 * (1)\] When the exponent is negative you flip it \[6/y^−5 = 6y^5\]

OpenStudy (anonymous):

Just to clarify here's a example\[y^-2 = 1/y^2\]